OpenStudy (anonymous):

For each function determine its domain and range: 1) f(x) = sqrt of x-4 2) srt of x^2-4 3) sqrt of 4-x^2 4) sqrt of x^2+44

OpenStudy (anonymous):

Sorry 2) shud say g(x) =...; 3) h(x)=... ; 4) ) k(x)=...

OpenStudy (anonymous):

Square root can't be negative! Just figure out which numbers make those negative and it can't be! I'll do 1 for you. x-4 has to be greater than or equal to 0 \[x - 4 \ge 0\] \[x \ge 4 \] is the domain the range (y values) can never be negative (we just said that!) so \[y \ge 0 \]

OpenStudy (anonymous):

\[f(x) = \sqrt{x-4}\].....this is how number 1 looks

OpenStudy (anonymous):

i see... so the domains and ranges are teh numbers wil make teh sqaure root positive

OpenStudy (anonymous):

is that correct? and then i just write an inequality for that?

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

@chaise...are you there?